How many oxygen atoms are in 1.60 g

g of quartz?

number of O atoms =



1
Expert's answer
2021-05-19T02:28:20-0400

M(SiO2) = 60.08 g/mol

n(SiO2) "= \frac{m}{M} = \frac{1.60}{60.08}=0.0266 \;mol"

Number of molecules:

N(SiO2) "= 0.0266 \times 6.02 \times 10^{23}=1.603 \times 10^{22}"

Each SiO2 molecule contains 2 oxygen atoms

number of O atoms "= 2 \times 1.603 \times 10^{22} = 3.206 \times 10^{22}"